# Interpreting Context Look-ups in Transformers: Investigating Attention-MLP Interactions

### v1

@@ Abstract
In this paper, we investigate the interplay between attention heads and specialized “next-token” neurons in the Multilayer Perceptron that predict specific tokens. By prompting an LLM like GPT-4 to explain these model internals, we can elucidate attention mechanisms that activate certain next-token neurons. Our analysis identifies attention heads that recognize contexts relevant to predicting a particular token, activating the associated neuron through the residual connection. We focus specifically on heads in earlier layers consistently activating the same next-token neuron across similar prompts. Exploring these differential activation patterns reveals that heads that specialize for distinct linguistic contexts are tied to generating certain tokens. Overall, our method combines neural explanations and probing isolated components to illuminate how attention enables context-dependent, specialized processing in LLMs.
@@
Clement ${}^{\ast\heartsuit\spadesuit}$ Shay B. ${}^{\diamondsuit}$  Fazl ${}^{\ast\heartsuit\clubsuit}$$^{\heartsuit}$part Research ${}^{\diamondsuit}$chool of Informatics, University of Edinburgh ${}^{\clubsuit}$epartment of Engineering Sciences, University of Oxford ${}^{\spadesuit}$chool of Computer Science and Engineering, Nanyang Technological University
@@ 1 Introduction
Figure 1: Our approach for characterizing attention heads. For a given token-predicting neuron, we run multiple prompts through GPT-2 Large to find attention heads that activate the neuron. We find that some attention heads activate the neuron only in specific contexts, and use GPT-4 to automate this discovery.
Recent advancements in natural language understanding and generation have been primarily dominated by Transformer-based models. GPT-4 OpenAI (2023) stands out as a particularly impressive achievement in this domain. Despite the accurate and sophisticated text generation capabilities of these models, a clear understanding of the inner workings of large language models (LLMs) remains elusive. This gap in knowledge further obscures the mechanistic functional forms of LLMs when their behavior is contrary to the common belief Miceli Barone et al. (2023), or when it is challenging to modify their factual knowledge Hoelscher-Obermaier et al. (2023). Gaining a comprehensive understanding of these models is crucial for expanding their theoretical foundations, improving real-world applications, mitigating biases, and reducing unsafe and potentially catastrophic risks. In this line, we aim to provide insight into the mechanistic interpretability of GPT-2 (Radford et al., 2017) by examining the associations between its attention heads and its multi-layer perceptrons (MLPs) neurons.
Previous research has focused on attention mechanisms, specifically multi-head self-attention and MLPs, to model complex input dependencies and generate text. Separate analyses of attention heads and MLPs have shown that these components enable the models to tackle a wide range of language understanding and generation tasks (Ferrando et al., 2023). However, the exact roles, combinations, and interactions of the attention heads and MLPs have yet to be fully elucidated. This is especially true in large-scale transformers like GPT-3 Brown et al. (2020) and GPT-4 OpenAI (2023). We aim to characterize attention heads and their interaction with next-token neurons, exploring their dynamics and shedding light on the transformer’s generative capabilities. To achieve this, we propose a hierarchical framework to systematically process and refine model inputs (prompts). This method allows us to determine the most influential attention heads for a next-token neuron for a given prompt. Furthermore, it enables us to quantify the individual attention head’s contribution to the language understanding and generation process through its head attribution scores and head explanation scores. Through this framework, we investigate the complex interplay between attention heads and next-token neurons, gaining a better understanding of their cooperative roles within GPT-2 Radford et al. (2019).
By conducting this in-depth investigation, we aim to uncover insights that can guide future advancements in transformer-based language modeling, facilitating enhanced prediction, interpretation, mitigating unsafe behavior and control. Additionally, this understanding proves crucial for identifying potential biases and inaccuracies within the model, promoting more responsible and immediate reliable real-life deployment of language model backed systems.
@@
Our main contributions are as follows: (1) We find that attention heads can recognize contexts relevant to predicting a token, and activate a downstream token-predicting neuron to help predict that token associated with that context; (2) We develop an automated analysis method to explain and characterize the interactions between attention heads and token-predicting neurons; (3) We propose an automated approach to evaluate the quality and effectiveness of these explanations.
@@ 2.2 MLP Interpretability
The final approach focuses on the analysis of individual neurons by inspecting their activation patterns or weights. Bills et al. (2023) and Foote et al. (2023) found that certain neurons consistently activate for specific concepts or syntax patterns. To study interactions between attention layers and MLP layers, it would be simpler to consider a neuron as a single analytical unit.
@@
In particular, Geva et al. (2022) analyzed neuron activation patterns in vocabulary space and identified neurons tied to predicting particular tokens. Building on this, Miller and Neo (2023) isolated a neuron specialized for the token "an" based on its output weights and prediction correlation. We term these types of neurons "next-token neurons."
@@ 2.3 Automated Interpretability
In particular, this approach allows us to explore the interaction between attention heads and next-token neurons, shedding light on how attention focuses processing to predict certain tokens. Our goal is thus to use GPT-4’s explanations to better understand attention head behavior in activating specific next-token neurons.
@@ 3 Background
Transformer Models. We examine transformer-based Language Models (LLMs) characterized by a sequence of length $L$ and dimension $d$, represented as $\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\in\mathbb{R}^{d}$. Each attention head is defined by parameter matrices $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$. Specifically, for token $i$, we compute $k_{i}=W_{k}\mathbf{x}_{i}$, $q_{i}=W_{q}\mathbf{x}_{i}$, and $v_{i}=W_{v}\mathbf{x}_{i}$. Furthermore, we define $K_{i}=[k_{1},\ldots,k_{i}]\in\mathbb{R}^{d\times i}$ and $V_{i}=[v_{1},\ldots,v_{i}]\in\mathbb{R}^{d\times i}$. The output of the head at token $i$ is denoted by $\mathbf{o}_{i}\in\mathbb{R}^{d}$, computed as $\mathbf{o}_{i}=V_{i}\cdot\text{softmax}(K_{i}\cdot q_{i})$.
Attention Head. An embedding, represented by function $\Psi:\mathcal{D}\rightarrow\mathbb{R}^{d}$, maps elements from domain $\mathcal{D}$ to vectors in $\mathbb{R}^{d}$. Meanwhile, a Multi-Layer Perceptron (MLP) is a neural network layer that processes vectors of size $dl$, transforming them into vectors of the same size. This transformation is achieved through a function $f:\mathbb{R}^{dl}\rightarrow\mathbb{R}^{dl}$, parameterized by weight matrices $W_{1}$ and $W_{2}$, and involves applying an activation function $\sigma$. Both the embedding and the MLP layer operate individually on each token (Vaswani et al., 2017).
@@ 3.1 Relating Attention Heads and Next-Token Neurons
We can use this interaction to analyze the activity of an attention head in relation to a next-token neuron. For example, if an attention head is highly responsible for activating the “go” neuron more frequently in certain prompts compared to others, we can look at the two sets of prompts to find out whether the attention head is responding to specific contexts. In our methodology, we study GPT-2 Large (774M parameters), and use GPT-4 to automate this analysis in an approach similar to Bills et al. (2023).
@@ 4.2 High-Activating Prompts
To study these neurons in their max-activating contexts, we define a function $\psi(p)$ that finds the max activation a neuron achieves for any token position within a prompt $p$. To identify these high-activation prompts, we initially run 10,000 prompts through the model. From these, we select the top 20 prompts for each neuron, noted as $P=[p_{1},p_{2},...,p_{30}]$, which is sufficient for finding interpretable activity in their associated attention heads. We further discuss the number of prompts used in §5.2.1.
After identifying these top 20 prompts, we refine them for clarity and conciseness by truncating the prompts from the beginning as long as they retain 80% of their initial activation, represented by the condition $\psi(p_{i}^{*})\geq 0.8\cdot\psi(p_{i})$ where $p_{i}^{*}$ is the shortened prompt. We then discard any prompts exceeding 100 tokens in length and replace line breaks with spaces. The method and its impact on prompt clarity is shown in in Figure 4.
@@
Figure 4: An example of prompt truncation. The activation of the " go" neuron is a minimum of 80% as compared to its activation for the original prompt.
The 10,000 prompts are from The Pile (Gao et al., 2020) [^1] Released under the MIT license., a diverse, open-source dataset composed of several smaller datasets of different domains, including code, books, and academic writing. By using the Pile, we can identify prompts in which next-token neurons activate over different contexts. This helps to ensure that the model behavior is sufficiently general across different contexts. The dataset is further discussed in §5.5.
@@ 4.4 Explaining Attention Heads
given the counts of true positives TP, false positives FP, true negatives TN, and false negatives FN. If there is no classification made for one class, and either the true positive rate or true negative rate is not applicable, we assign the score to be the other. We also remove cases where all test prompts fall under one classes, as classification can succeed trivially.
@@ 5.1 Attention Heads May Capture Phrases or Context
Our analysis reveals that certain attention heads in GPT-4 exhibit consistent activity patterns for prompts that maximally activate specific next-token neurons. For instance, we find heads that activate the "as"-predicting neuron only in contexts like "as early… as" or "as high… as", but not other prompts without these phrases (see Table 1). When clear differences exist between activating and non-activating prompts, GPT-4 can often generate good explanations of the head’s specialized function.
@@
Several heads have high "head explanation scores" (Figure 5), meaning they reliably activate for specific phrases or word usages. Different heads capture various uses of the same word that activate its predicting neuron. For example, the "number" neuron has one head that activates it for ranking concepts (like "number one") and another head for identification numbers ("mobile number", "passport number"). Table 2 and Appendix E provides examples of explainable heads.
@@
($31,3621,538$) "only" Describes a specific condition, purpose, or limitation
@@
Table 2: Attention Head-Neuron pairs and the explanation for the activity of the attention head. Attention heads can activate token-predicting neurons for a specific context or phrase.
@@
(f) Pythia 160M)
@@ 5.3 Finding Explainable Heads in Other Models
The head explanation score distributions for these models skewed rightwards (Figures 5). These rightward skews indicate that the head explanations provided useful information to classify prompts above random chance.
@@ 5.4 Verification by Head Ablation
To validate the faithfulness of the neural explanations, we ablate selected attention heads when running test prompts through the model. If the explanations fully capture head behavior, the token probability should decrease only on prompts where the head is described as active. We perform ablation experiments by zeroing out individual heads and comparing token probabilities between original and ablated models on the same prompts. As Figure 6 shows, token probabilities decreased significantly more on prompts labeled as head-active versus head-inactive.
@@ 6 Conclusion
We investigated attention heads and their role in activating next-token neurons. We found that some attention heads would only activate their associated next-token neurons in a subset of max-activating prompts. We then use GPT-4 to generate explanations for this activity, and made GPT-4 use that explanation for a classification task on a test-set of prompts to evaluate explanation quality. Our findings reveal that attention heads process context and use next-token neurons to help predict a word associated with, but not previously seen in that context. This provides insight into how attention heads can work with MLP neurons to predict new words. Future work could investigate whether this phenomenon happens to such neurons throughout their range of activations, which could give insights to the role of next-token neurons when the predicted token is unlikely to be the neuron’s associated token. Further research could also look into whether multiple attention heads could work with a single next-token neuron in a similar manner, or whether a single attention head could work with multiple next-token neurons in a similar manner. This could provide insights to whether such attention-head neuron interactions extend to more complex contexts. In conclusion, our study provided a strong insights into GPT-4’s generation and understanding process by examining head attribution scores and explanation scores between attention heads and neurons.
@@
Our investigation relied on a maximum of 30 prompts per neuron. We also observed that GPT-4 struggles to generate coherent and accurate explanations in response to longer prompts, influencing the quality and consistency of explanations provided. To address these problems, we could feed more prompts per neuron into GPT-4, and perform few-shot explanation generation instead of zero-shot. Notably, Bills et al. (2023) used few-shot generation to explain neuron activity. However, this would incur higher computational costs with GPT-4, and it is necessary to strike a balance between thorough research and resource constraints.
Lastly, our scope of our analysis may not completely capture complex attention head and neuron behavior. We scope only encompassed neurons whose associated tokens are whole words. Our analysis is a starting point to characterizing such behavior. Therefore, we encourage in future research to employ additional datasets, benchmarking tasks, and examination techniques to enhance our understanding of LLMs strengths and limitations regarding neuron activation, explanation generation, and various aspects of its learned knowledge.

### v2

@@ Abstract
Understanding the inner workings of large language models (LLMs) is crucial for advancing their theoretical foundations and real-world applications. While the attention mechanism and multi-layer perceptrons (MLPs) have been studied independently, their interactions remain largely unexplored. This study investigates how attention heads and next-token neurons interact in LLMs to predict new words. We propose a methodology to identify next-token neurons, find prompts that highly activate them, and determine the upstream attention heads responsible. We then generate and evaluate explanations for the activity of these attention heads in an automated manner. Our findings reveal that some attention heads recognize specific contexts relevant to predicting a token and activate a downstream token-predicting neuron accordingly. This mechanism provides a deeper understanding of how attention heads work with MLP neurons to perform next-token prediction. Our approach offers a foundation for further research into the intricate workings of LLMs and their impact on text generation and understanding.
@@
Clement Neo∗♡♠ Shay B. Cohen  Fazl Barez∗♡♣ Apart Research Nanyang Technological University School of Informatics, University of Edinburgh Department of Engineering Sciences, University of Oxford
@@ 1 Introduction
Figure 1: Our approach for characterizing attention heads. For a given token-predicting neuron, we run multiple prompts through GPT-2 and Pythia to find attention heads that activate the neuron. We find that some attention heads activate the neuron only in specific contexts, and use GPT-4 to automate this discovery.
Transformer-based models have made significant advancements in natural language understanding and generation. However, the inner workings of large language models (LLMs) remain largely unexplored, particularly when their behavior deviates from common beliefs Miceli Barone et al. (2023) or when modifying their factual knowledge proves challenging Hoelscher-Obermaier et al. (2023). A comprehensive understanding of these models is essential for expanding their theoretical foundations, improving real-world applications, mitigating biases, and reducing potential risks.
This study aims to provide insight into the mechanistic interpretability of GPT-2 (Radford et al., 2019) and Pythia (Biderman et al., 2023) by examining the associations between their attention heads and multi-layer perceptron (MLP) neurons. Previous research has analyzed attention mechanisms and MLPs separately, demonstrating their ability to model complex input dependencies and generate text (Ferrando et al., 2023). However, the exact roles, combinations, and interactions of attention heads and MLPs require further investigation.
To address this, we propose a methodology to investigate the interactions between attention heads and next-token neurons. We first identify next-token neurons and find prompts that highly activate them. We then determine the attention heads most responsible for activating each next-token neuron during the forward pass, using a head attribution score. Next, we use GPT-4 to generate explanations for the activity patterns of these attention heads. Finally, we evaluate the explanations’ quality by using GPT-4 as a zero-shot classifier to predict head activity on new prompts, comparing the predictions to actual head activity.
Our main contributions are as follows:
  1. 1.
@@
    We find that attention heads can recognize contexts relevant to predicting a token, and activate a downstream token-predicting neuron to help predict that token associated with that context;
  2. 2.
    We develop an automated analysis method to explain and characterize the interactions between attention heads and token-predicting neurons;
  3. 3.
    We propose an automated approach to evaluate the quality and effectiveness of these explanations.
@@ 2.2 MLP Interpretability
A more recent approach focuses on the analysis of individual neurons by inspecting their activation patterns or weights. Bills et al. (2023) and Foote et al. (2023) found that certain neurons consistently activate for specific concepts or syntax patterns. To study interactions between attention layers and MLP layers, it would be simpler to consider a neuron as a single analytical unit.
@@
In particular, Geva et al. (2022) analyzed neuron activation patterns in vocabulary space and identified neurons tied to predicting particular tokens. Building on this, Miller and Neo (2023) isolated a neuron specialized for the token “an” based on its output weights and prediction correlation. We term these types of neurons “next-token neurons”.
@@ 2.3 Automated Interpretability
In particular, this approach allows us to examine the interaction between attention heads and next-token neurons, shedding light on how attention focuses processing to predict certain tokens. Our goal is thus to use GPT-4’s explanations to better understand attention head behavior in activating specific next-token neurons.
@@ 3 Background
Transformer Models. We examine transformer-based Language Models (LLMs) characterized by a sequence of length $L$ and dimension $d$, represented as $\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\in\mathbb{R}^{d}$. Each attention head is defined by parameter matrices $W_{k},W_{q},W_{v}\in\mathbb{R}^{d\times d}$. Specifically, for token $i$, we compute $k_{i}=W_{k}\mathbf{x}_{i}$, $q_{i}=W_{q}\mathbf{x}_{i}$, and $v_{i}=W_{v}\mathbf{x}_{i}$. Furthermore, we define $K_{i}=[k_{1},\ldots,k_{i}]\in\mathbb{R}^{d\times i}$ and $V_{i}=[v_{1},\ldots,v_{i}]\in\mathbb{R}^{d\times i}$. The output of the head at token $i$ is denoted by $\mathbf{o}_{i}\in\mathbb{R}^{d}$, computed as $\mathbf{o}_{i}=V_{i}\cdot\text{softmax}(\frac{K_{i}\cdot q_{i}}{\sqrt{d_{k}}})$.
An embedding, represented by function $\Psi:\mathcal{D}\rightarrow\mathbb{R}^{d}$, maps elements from domain $\mathcal{D}$ to vectors in $\mathbb{R}^{d}$. Meanwhile, a Multi-Layer Perceptron (MLP) is a neural network layer that processes vectors of size $dl$, transforming them into vectors of the same size. This transformation is achieved through a function $f:\mathbb{R}^{dl}\rightarrow\mathbb{R}^{dl}$, parameterized by weight matrices $W_{1}$ and $W_{2}$, and involves applying an activation function $\sigma$. Both the embedding and the MLP layer operate individually on each token (Vaswani et al., 2017).
@@ 3.1 Relating Attention Heads and Next-Token Neurons
We can use this interaction to analyze the activity of an attention head in relation to a next-token neuron. For example, if an attention head is highly responsible for activating the “go” neuron more frequently in certain prompts compared to others, we can look at the two sets of prompts to find out whether the attention head is responding to specific contexts. In our methodology, we study several models including GPT-2 Large (774M parameters), and use GPT-4 to automate this analysis in an approach similar to Bills et al. (2023).
@@ 4.2 High-Activating Prompts
To study these neurons in their max-activating contexts, we define a function $\psi(p)$ that finds the max activation a neuron achieves for any token position within a prompt $p$. To identify these high-activation prompts, we initially run 10,000 prompts through the model. From these, we select the top 20 prompts for each neuron, noted as $P=[p_{1},p_{2},...,p_{20}]$, which is sufficient for finding interpretable activity in their associated attention heads. We further discuss the number of prompts used in §5.2.1.
Prompt pre-processing. After identifying these top 20 prompts, we refine them for clarity and conciseness by truncating the prompts from the beginning to the point where they retain 80% of their initial activation, represented by the condition $\psi(p_{i}^{*})\geq 0.8\cdot\psi(p_{i})$ where $p_{i}^{*}$ is the shortened prompt. We then discard any prompts exceeding 100 tokens in length and replace line breaks with spaces. The method and its impact on prompt clarity is shown in in Figure 4.
@@
Figure 4: An example of prompt truncation. The activation of the " go" neuron is a minimum of 80% as compared to its activation for the original prompt. This pre-processing step automatically shortens the prompt while still making sure it activates the neuron significantly.
Dataset source. The 10,000 prompts are from The Pile (Gao et al., 2020) [^1] Released under the MIT license., a diverse, open-source dataset composed of several smaller datasets of different domains, including code, books, and academic writing. By using the Pile, we can identify prompts in which next-token neurons activate over different contexts. This helps to ensure that the model behavior is sufficiently general across different contexts. The dataset is further discussed in §5.5.
@@ 4.4 Explaining Attention Heads
given the counts of true positives TP, false positives FP, true negatives TN, and false negatives FN. These counts are derived from comparing GPT-4’s classifications against the ground truth of head activity in our test set of 10 prompts, where positive examples are prompts with active heads and negative examples are those with inactive heads. If there is no classification made for one class, and either the true positive rate or true negative rate is not applicable, we assign the score to be the other. We also remove cases where all test prompts fall under one classes, as classification can succeed trivially.
@@ 5.1 Attention Heads May Capture Phrases or Context
Our analysis reveals that certain attention heads in GPT-2 exhibit consistent activity patterns for prompts that maximally activate specific next-token neurons. For instance, we find heads that activate the "as"-predicting neuron only in contexts like "as early… as" or "as high… as", but not other prompts without these phrases (see Table 1). When clear differences exist between activating and non-activating prompts, GPT-4 can often generate good explanations of the head’s specialized function.
@@
Several heads have high "head explanation scores" (Figure 5), meaning they reliably activate for specific phrases or word usages. Different heads capture various uses of the same word that activate its predicting neuron. For example, the "number" neuron has one head that activates it for ranking concepts (like "number one") and another head for identification numbers ("mobile number", "passport number"). Table 2 and Appendices E, F provides examples of explainable heads.
@@
$(31,3621,538)$only" Describes a specific condition, purpose, or limitation
@@
Table 2: This table illustrates how specific attention heads activate token-predicting neurons in certain contexts. The notation (Layer, Neuron, Head) is used, where ’Head’ is a flattened index. For example, (31, 3621, 538) refers to the 3621st neuron in layer 31, paired with the attention head corresponding to flattened index 538. More examples in Appendix F.
@@
$\text{Mean}=0.575$
KS $\text{KS }p\text{-value}=3.2\times 10^{-12}$
@@
$\text{Mean}=0.610$
KS $\text{KS }p\text{-value}=4.13\times 10^{-13}$
@@
$\text{Mean}=0.650$
KS $\text{KS }p\text{-value}=1.6\times 10^{-14}$
@@
$\text{Mean}=0.574$
KS $\text{KS }p\text{-value}=8.6\times 10^{-7}$
(f) Pythia 160M
@@
$\text{Mean}=0.621$
KS $\text{KS }p\text{-value}=1.6\times 10^{-7}$
@@ 5.3 Finding Explainable Heads in Other Models
The head explanation score distributions for these models skewed rightwards (Figures 5). These rightward skews indicate that the head explanations provided useful information to classify prompts above random chance. However, the presence of skew alone is not sufficient to draw this conclusion.
To further support this finding, we also find that the mean head explanation scores for all models were higher than the random baseline. Additionally, the Kolmogorov-Smirnov test revealed a a statistically significant difference between the score distributions of the model and the random baseline (values in Figure 5). These results strengthen the evidence that the head explanations offer valuable information for prompt classification beyond random chance.
@@ 5.4 Verification by Head Ablation
To validate the faithfulness of the neural explanations, we ablate selected attention heads when running test prompts through the model. If the explanations fully capture head behavior, the token probability should decrease only on prompts where the head is described as active. We perform ablation experiments by zeroing out individual heads and comparing token probabilities between original and ablated models on the same prompts. As Figure 6 shows, there is a statistically significant decrease in token probabilities on prompts labeled as head-active versus head-inactive.
@@ 6 Conclusion
We investigated attention heads and their role in activating next-token neurons. We found that some attention heads would only activate their associated next-token neurons in a subset of max-activating prompts. We then use GPT-4 to generate explanations for this activity, and made GPT-4 use that explanation for a classification task on a test-set of prompts to evaluate explanation quality. Our findings reveal that attention heads process context and use next-token neurons to help predict a word associated with, but not previously seen in that context. This provides insight into how attention heads can work with MLP neurons to predict new words.
Future work could investigate whether this phenomenon happens to such neurons throughout their range of activations, which could give insights to the role of next-token neurons when the predicted token is unlikely to be the neuron’s associated token. Further research could also look into whether multiple attention heads could work with a single next-token neuron in a similar manner, or whether a single attention head could work with multiple next-token neurons in a similar manner. This could provide insights to whether such attention-head neuron interactions extend to more complex contexts.
@@
Our investigation relied on a maximum of 30 prompts per neuron. We also observed that GPT-4 struggles to generate coherent and accurate explanations in response to longer prompts, influencing the quality and consistency of explanations provided. To address these problems, we could feed more prompts per neuron into GPT-4, and perform few-shot explanation generation instead of zero-shot. Notably, Bills et al. (2023) used few-shot generation to explain neuron activity. However, this would incur higher financial costs with GPT-4, and it is necessary to strike a balance between thorough research and resource constraints.
Lastly, the scope of our analysis may not completely capture complex attention head and neuron behavior. We scope only encompassed neurons whose associated tokens are whole words, and we limit our analysis to prompts that are shorter than 100 tokens. Our analysis is a starting point to characterizing such behavior. Therefore, we encourage in future research to employ additional datasets, benchmarking tasks, and examination techniques to enhance our understanding of LLMs strengths and limitations regarding neuron activation, explanation generation, and various aspects of its learned knowledge.
Acknowledgements
We thank Ameya Prabhu, Yftah Ziser, Jason Schreiber for reading drafts of the paper.